write the standard form of the equation of the line that passes through (5,-2)
and is parallel to the line with equation 3x+2y+4=0

Answer by scott8148(6628)

parallel means the same slope

so the x and y coefficients (A and B) are the same

substitute the coordinates of the point into the equation to find a new value for the constant term (C)

3(5) + 2(-2) + C = 0 ___ 15 - 4 + C = 0 ___ C = -11

the new equation is ___ 3x + 2y - 11 = 0